# Computational Complexity of Physical Information Sufficiency

### v3

@@ A.7 Complete Claim Coverage Matrix
thm:cost-asymmetry-eth Full DQ.ClaimClosure.cost_asymmetry_eth_conditional, DQ.HardnessDistribution.isWrongHardness
thm:dichotomy Full DQ.ClaimClosure.dichotomy_conditional, DQ.ClaimClosure.explicit_state_upper_core, DQ.ClaimClosure.hard_family_all_coords_core
thm:exact-certified-gap-iff-admissible Full DQ.ClaimClosure.epsilon_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core, DQ.ClaimClosure.exact_raw_eq_certified_iff_certainty_inflation_core, DQ.Physics.PhysicalCore.admissible_report_requires_strict_certified_work
thm:generalized-dominance Full DQ.HardnessDistribution.generalized_dominance_can_fail_without_wrong_growth
@@
thm:generalized-saturation-possible Full DQ.HardnessDistribution.gap_conservation_card, DQ.HardnessDistribution.right_dominates_wrong
@@
thm:linear-saturation-iff-zero Full DQ.HardnessDistribution.simplicityTax_grows
thm:overmodel-diagnostic Full DQ.ClaimClosure.boundaryCharacterized_iff_exists_sufficient_subset, DQ.ClaimClosure.no_exact_identifier_implies_not_boundary_characterized
@@
thm:physical-bridge-bundle Full CT.physical_bridge_bundle
thm:regime-coverage Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.regime_core_claim_proved
thm:tax-conservation Full DQ.HardnessDistribution.complete_model_dominates_after_threshold
thm:tax-grows Full DQ.HardnessDistribution.totalDOF_ge_intrinsic
thm:tractable Full DQ.ClaimClosure.thermo_mandatory_cost_core, DQ.ClaimClosure.tractable_bounded_core, DQ.ClaimClosure.tractable_separable_core, DQ.ClaimClosure.tractable_subcases_conditional
@@
thm:typed-completeness-static Full DQ.ClaimClosure.typed_claim_admissibility_core
Auto summary: mapped 123/123 (full=123, derived=0, unmapped=0).
@@ A.8 Proof Hardness Index (Auto)
cor:gap-externalization cost-growth H=cost-growth HD18, HD21
cor:generalized-eventual-dominance cost-growth H=cost-growth HD9
cor:linear-positive-no-saturation cost-growth H=cost-growth HD15
cor:right-wrong-hardness cost-growth H=cost-growth HD16
cor:type-system-threshold cost-growth H=cost-growth HD14
@@
prop:generalized-assumption-boundary cost-growth H=cost-growth HD6, HD7
prop:hardness-conservation cost-growth H=cost-growth HD20
prop:hardness-efficiency-interpretation cost-growth H=cost-growth HD10
@@
thm:generalized-dominance cost-growth H=cost-growth HD8
thm:generalized-saturation-possible cost-growth H=cost-growth HD5, HD17
@@
cor:finite-budget-no-exact-admissibility exp-lb-conditional H=exp-lb-conditional DQ573, DQ574
cor:finite-budget-threshold-impossibility exp-lb-conditional H=exp-lb-conditional DQ574, DQ568, DQ569
cor:no-universal-survivor-no-succinct-bound exp-lb-conditional H=exp-lb-conditional DQ578
@@
prop:eventual-explicit-infeasibility exp-lb-conditional H=exp-lb-conditional DQ577
prop:least-divergence-point exp-lb-conditional H=exp-lb-conditional DQ576
prop:payoff-threshold exp-lb-conditional H=exp-lb-conditional DQ575, DQ579
prop:policy-closure-beyond-divergence exp-lb-conditional H=exp-lb-conditional DQ580, DQ581
@@
cor:query-obstruction-bool query-lb H=query-lb CC49, L6
prop:checking-witnessing-duality query-lb H=query-lb DQ587, DQ588
prop:query-finite-state-generalization query-lb H=query-lb L7, L13
prop:query-randomized-robustness query-lb H=query-lb L5, L11, L12
prop:query-randomized-weighted query-lb H=query-lb L23, L24
prop:query-tightness-full-scan query-lb H=query-lb L10
prop:query-weighted-transfer query-lb H=query-lb L21, L22
@@
cor:practice-bounded tractable-structured H=tractable-structured L15
cor:practice-structured tractable-structured H=tractable-structured L16
cor:practice-tree tractable-structured H=tractable-structured L17
@@
cor:exact-no-competence-zero-certified unspecified AR,DC IC41
@@
cor:hardness-raw-only-exact unspecified AR CC21, IC23
@@
cor:physical-counterexample-core-failure unspecified AR,E,Qb,Qf,S,S+ETH CT3, CT7, CT8, DQ571
cor:physics-no-universal-exact-claim unspecified AR,CR,DC CC38
cor:practice-unstructured unspecified - L1
cor:rlff-abstain-no-certs unspecified AR IC30, DQ570
prop:abstain-guess-self-signal unspecified AR,DC IC46
@@
prop:certified-confidence-gate unspecified AR,DC IC38, IC39
prop:decision-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ535, DQ537, DQ545, DQ546
@@
prop:empty-sufficient-constant unspecified AR,DC,DM DP6
prop:evidence-admissibility-equivalence unspecified AR,DC IC17, IC20, IC21
prop:exact-requires-evidence unspecified AR,DC IC35, IC36
prop:fraction-defined-under-bound unspecified AR,DC IC45
prop:heisenberg-strong-nontrivial-opt unspecified AR DQ584, DQ585, DQ586
prop:insufficiency-counterexample unspecified AR,DC,DM DP7, DP8
@@
prop:integrity-resource-bound unspecified AR,DC,S,S+ETH CC30, IC24, IC26
@@
prop:no-evidence-zero-certified unspecified AR,DC IC40
prop:no-fraction-without-bound unspecified AR,DC IC44
@@
prop:oracle-lattice-strict unspecified AR,E,Qb,Qf,S,S+ETH L3, L4
prop:oracle-lattice-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC43, L20
@@
prop:physical-claim-transport unspecified AR,E,LC,Qb,Qf,S,S+ETH,TA CT5, CT6, DQ572
prop:physics-no-universal-exact unspecified AR,CR,DC CC15
prop:query-regime-obstruction unspecified AR,E,Q_fin,Qb,Qf,S,S+ETH CC50, L7, L13
prop:query-state-batch-lb unspecified AR,E,Qb,Qf,S,S+ETH L8, L14
prop:query-subproblem-transfer unspecified AR,E,Qb,Qf,S,S+ETH CC52, CC58, CC59
prop:query-value-entry-lb unspecified AR,E,Qb,Qf,S,S+ETH L9, L19
prop:raw-certified-bit-split unspecified AR CC11, IC10, IC11, IC13, IC14, IC15, IC16, DQ567
@@
prop:self-confidence-not-certification unspecified AR,DC IC47
@@
prop:snapshot-process-typing unspecified RA CC3, CC48, CC56
prop:steps-run-scalar unspecified AR,DC IC42, IC43
prop:substrate-unit-time unspecified AR,E,Qb,Qf,S,S+ETH DQ541, DQ542, DQ549
prop:time-discrete unspecified AR,E,Qb,Qf,S,S+ETH DQ547, DQ548
@@
prop:under-resolution-collision unspecified DM DQ582, DQ583
@@
thm:amortization unspecified LC HD3
thm:bridge-boundary-represented unspecified - CC8, CC9, CC10
thm:centralization-dominance unspecified AR,LC HD1, HD2
thm:claim-integrity-meta unspecified AR,CR,DC CC22
thm:config-reduction unspecified AR,E,S,TA CR1
thm:cost-asymmetry-eth unspecified S+ETH CC12, HD13
thm:dichotomy unspecified AR,E,Q_fin,Qb,Qf,RG,S,S+ETH CC16, CC23, CC24
thm:exact-certified-gap-iff-admissible unspecified AR CC17, CC18, CC20, DQ566
@@
thm:regime-coverage unspecified CR,S CC14, CC51
thm:tax-conservation unspecified AR,LC HD4
thm:tax-grows unspecified LC HD21
thm:tractable unspecified AR,E,Qb,Qf,RG,S,S+ETH,STR,TA CC69, CC70, CC71, CC72
thm:typed-completeness-static unspecified AR,CR,S CC76
Auto summary: indexed 123 claims by hardness profile (cost-growth=16; exp-lb-conditional=11; query-lb=8; succinct-hard=6; tractable-structured=6; unspecified=76).

### v4

@@ A.7 Complete Claim Coverage Matrix
thm:conservation Full DQ.InteriorVerification.agreeOnSet
thm:cost-asymmetry-eth Full DQ.ClaimClosure.bridge_transfer_iff_one_step_class, DQ.HardnessDistribution.linear_lt_exponential_plus_constant_eventually
thm:counting-gap Full DQ.Physics.BoundedAcquisition.counting_gap_theorem
thm:counting-gap-intro Full DQ.Physics.BoundedAcquisition.counting_gap_theorem
thm:cycle-cost Full DQ.ClaimClosure.RegimeSimulation, DQ.ClaimClosure.adq_ordering, DQ.ClaimClosure.system_transfer_licensed_iff_snapshot
thm:deficit-source Full DQ.InteriorVerification.interiorParetoDominates, DQ.InteriorVerification.interior_certificate_implies_non_rejection
thm:dichotomy Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.exact_raw_only_of_no_exact_admissible_core, DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
thm:discrete-acquisition Full DQ.Physics.BoundedAcquisition.acquisitions_are_transitions
thm:dq-physical Full DQ.ClaimClosure.DQ1, DQ.ClaimClosure.DQ2, DQ.ClaimClosure.DQ3, DQ.ClaimClosure.DQ4, DQ.ClaimClosure.DQ5, DQ.ClaimClosure.DQ6
thm:ec2-derived Full DQ.ClaimClosure.IA11, DQ.ClaimClosure.IA12, DQ.ClaimClosure.IA13, DQ.ClaimClosure.IA7, DQ.ClaimClosure.IA9
thm:ec3-derived Full DQ.Physics.LocalityPhysics.FPT10_ec3_is_logical, DQ.Physics.LocalityPhysics.FPT4_step_requires_distinct_moments, DQ.Physics.LocalityPhysics.FPT5_distinct_moments_positive_duration, DQ.Physics.LocalityPhysics.FPT6_step_takes_positive_time, DQ.Physics.LocalityPhysics.FPT8_propagation_takes_time
thm:energy-information Full DQ.ThermodynamicLift.energy_ge_kbt_nat_entropy
thm:entropy-rank Full DQ.numOptClasses_le_pow_srank_binary, DQ.quotientEntropy_le_srank_binary
thm:exact-certified-gap-iff-admissible Full DQ.ClaimClosure.declared_physics_no_universal_exact_certifier_core, DQ.ClaimClosure.dichotomy_conditional, DQ.ClaimClosure.exact_admissible_iff_raw_lt_certified_total_core, DQ.IntegrityCompetence.exact_raw_only_of_no_exact_admissible
thm:fi-coincide Full DQ.FunctionalInformation.first_principles_thermo_coincide, DQ.FunctionalInformation.functional_information_from_counting, DQ.FunctionalInformation.functional_information_from_thermodynamics
thm:fisher-rank-srank Full DQ.Statistics.fisherMatrix_rank_eq_srank, DQ.Statistics.sum_fisherScore_eq_srank
thm:gap-physical Full DQ.ClaimClosure.IA4
@@
thm:generalized-dominance Full DQ.HardnessDistribution.generalized_right_dominates_wrong_of_bounded_vs_identity_lower
thm:generalized-saturation-possible Full DQ.HardnessDistribution.generalizedTotal_with_saturation_eventually_constant, DQ.HardnessDistribution.saturatingSiteCost_eventually_constant
thm:growth-time Full DQ.InteriorVerification.GoalClass, DQ.InteriorVerification.InteriorDominanceVerifiable, DQ.InteriorVerification.TautologicalSetIdentifiable
thm:information-gap Full DQ.ClaimClosure.IA3
thm:landauer-structure Full DQ.Physics.LocalityPhysics.landauer_structure
@@
thm:linear-saturation-iff-zero Full DQ.HardnessDistribution.right_dominates_wrong
thm:measure-prerequisite Full DQ.Physics.MeasureNecessity.quantitative_claim_has_measure, DQ.Physics.MeasureNecessity.quantitative_measure_is_logical_prerequisite, DQ.Physics.MeasureNecessity.stochastic_claim_has_probability_measure, DQ.Physics.MeasureNecessity.stochastic_probability_is_logical_prerequisite
thm:nontriviality-counting Full DQ.Physics.LocalityPhysics.constant_function_singleton_image, DQ.Physics.LocalityPhysics.equal_states_constant_function, DQ.Physics.LocalityPhysics.information_requires_nontriviality, DQ.Physics.LocalityPhysics.singleton_image_zero_entropy, DQ.Physics.LocalityPhysics.trivial_states_all_equal, DQ.Physics.LocalityPhysics.triviality_implies_no_information, DQ.Physics.LocalityPhysics.zero_entropy_no_information
thm:orbital-cost Full DQ.ClaimClosure.AtomicCircuitExports.AC3, DQ.ClaimClosure.AtomicCircuitExports.AC4
thm:orbital-transition Full DQ.ClaimClosure.AtomicCircuitExports.AC1, DQ.ClaimClosure.AtomicCircuitExports.AC5
thm:overmodel-diagnostic Full DQ.ClaimClosure.anchor_query_relation_false_iff, DQ.ClaimClosure.no_exact_claim_admissible_under_hardness_core
@@
thm:physical-bridge-bundle Full DQ.Physics.ClaimTransport.physical_counterexample_yields_core_counterexample
thm:physical-incompleteness Full DQ.Physics.PhysicalIncompleteness.no_surjective_instantiation_of_card_gap, DQ.Physics.PhysicalIncompleteness.physical_incompleteness_of_bounds, DQ.Physics.PhysicalIncompleteness.physical_incompleteness_of_card_gap
thm:quotient-universal Full DQ.DecisionProblem.quotientMap_preservesOpt, DQ.DecisionProblem.quotient_is_coarsest, DQ.DecisionProblem.quotient_represents_opt_equiv
thm:rate-distortion-bridge Full DQ.Information.compression_below_srank_fails, DQ.Information.equiv_preserves_decision, DQ.Information.rate_distortion_bridge, DQ.Information.rate_equals_srank, DQ.Information.rate_monotone, DQ.Information.rate_zero_distortion, DQ.Information.shannonEntropy_nonneg, DQ.Information.srank_bits_sufficient
thm:regime-coverage Full DQ.ClaimClosure.cost_asymmetry_eth_conditional, DQ.ClaimClosure.poseAnchorQuery
thm:resolution-sufficient Full DQ.Physics.BoundedAcquisition.resolution_reads_sufficient
thm:second-law-counting Full DQ.Physics.LocalityPhysics.atypical_states_rare, DQ.Physics.LocalityPhysics.entropy_is_information, DQ.Physics.LocalityPhysics.errors_accumulate, DQ.Physics.LocalityPhysics.random_misses_target, DQ.Physics.LocalityPhysics.second_law_from_counting, DQ.Physics.LocalityPhysics.verification_is_information, DQ.Physics.LocalityPhysics.wrong_paths_dominate
thm:six-subcases Full DQ.ClaimClosure.subproblem_hardness_lifts_to_full, DQ.ClaimClosure.subproblem_transfer_as_regime_simulation, DQ.ClaimClosure.sufficiency_conp_complete_conditional, DQ.ClaimClosure.sufficiency_conp_reduction_core, DQ.StochasticSequential.ClaimClosure.claim_tractable_subcases_to_P
thm:srank-physical Full DQ.ClaimClosure.SR1, DQ.Physics.BoundedAcquisition.energy_ge_srank_cost, DQ.Physics.BoundedAcquisition.srank_le_resolution_bits
thm:substrate-degradation Full DQ.ClaimClosure.SE1, DQ.ClaimClosure.SE2, DQ.ClaimClosure.SE3, DQ.ClaimClosure.SE4
thm:tax-conservation Full DQ.HardnessDistribution.gap_conservation_card
thm:tax-grows Full DQ.HardnessDistribution.totalExternalWork_eq_n_mul_gapCard
thm:thermo-derived Full DQ.Physics.BoundedAcquisition.physical_grounding_bundle
thm:topology-motion Full DQ.ClaimClosure.declaredRegimeFamily_complete, DQ.ClaimClosure.explicit_assumptions_required_of_not_excused_core
thm:tractable Full DQ.ClaimClosure.subproblem_hardness_lifts_to_full, DQ.ClaimClosure.subproblem_transfer_as_regime_simulation, DQ.ClaimClosure.sufficiency_conp_complete_conditional, DQ.ClaimClosure.sufficiency_conp_reduction_core
thm:tur-bridge Full DQ.Physics.multiple_futures_entropy_production, DQ.Physics.transitionProb_nonneg, DQ.Physics.transitionProb_sum_one, DQ.Physics.tur_bridge
@@
thm:typed-completeness-static Full DQ.ClaimClosure.thermo_conservation_additive_core
thm:universe-membership Unmapped (no derived Lean handle found)
thm:wasserstein-bridge Full DQ.Physics.integrity_is_centroid, DQ.Physics.single_future_zero_cost, DQ.Physics.transportCost_pos_of_offDiag, DQ.Physics.wasserstein_bridge
Notes: (1) Full rows come from theorem-local inline anchors in this paper. (2) Derived rows are filled by dependency/scaffold claim-handle derivation (same paper-handle label across proof dependencies). (3) Unmapped means no local anchor and no derivable dependency support were found.
Auto summary: mapped 220/222 (full=220, derived=0, unmapped=2).
@@ A.8 Proof Hardness Index (Auto)
cor:channel-degradation cost-growth H=cost-growth,Q_fin CH2, CH6
cor:finite-lifetime cost-growth H=cost-growth,Q_fin SE5
cor:gap-externalization cost-growth H=cost-growth HD21, HD26
cor:generalized-eventual-dominance cost-growth H=cost-growth HD10
cor:hardness-raw-only-exact cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC21, IC23
cor:linear-positive-no-saturation cost-growth H=cost-growth HD16
cor:no-free-computation cost-growth AR,H=cost-growth CC7, CC5, CC8
cor:physical-counterexample-core-failure cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CT6, CT5, CT4
cor:right-wrong-hardness cost-growth H=cost-growth HD19
cor:speed-integrity cost-growth H=cost-growth,Q_fin SE6
cor:thermo-dq cost-growth AR,H=cost-growth DQ7, DQ8
cor:type-system-threshold cost-growth H=cost-growth HD15
cor:zero-capacity cost-growth H=cost-growth,Q_fin CH3
prop:decision-unit-time cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT13, DT11, DT12, DT10
@@
prop:generalized-assumption-boundary cost-growth H=cost-growth HD7, HD8
prop:hardness-conservation cost-growth H=cost-growth HD23, HD25
prop:hardness-efficiency-interpretation cost-growth H=cost-growth HD11
prop:oracle-lattice-strict cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC27, CC29
prop:oracle-lattice-transfer cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC43, CC52
prop:query-regime-obstruction cost-growth AR,E,H=cost-growth,LC,Q_fin,Qb,Qf,RG,S,S+ETH CC50
prop:query-state-batch-lb cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC28, CC50
prop:query-subproblem-transfer cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC52, CC58, CC59
prop:query-value-entry-lb cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC29, CC50
prop:raw-certified-bit-split cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC11, IC10, IC11, IC13, IC14, IC15, IC16, IC18, IC19
prop:reaction-competence cost-growth H=cost-growth,Q_fin MI3, MI4
prop:run-time-accounting cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT18, DT19, DT16, DT15
prop:stochastic-landauer-floor cost-growth AR,H=cost-growth DC36, DC35, DC37
prop:structural-asymmetry cost-growth AR,H=cost-growth SR1, SR2, SR3
prop:substrate-unit-time cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT23, DT22, DT24
@@
prop:time-discrete cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH DT7, DT6
thm:amortization cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH HD4
thm:bayes-from-dq cost-growth AR,H=cost-growth FN14, BB4, BB5, BB3
thm:bridge-boundary-represented cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH CC8, CC9, CC10
thm:centralization-dominance cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH HD1, HD2
thm:choice-pays cost-growth AR,H=cost-growth GE3, GE7
thm:conservation cost-growth H=cost-growth,Q_fin IV4
thm:cycle-cost cost-growth AR,H=cost-growth CC1, CC2, CC3
thm:deficit-source cost-growth H=cost-growth,Q_fin IV5, IV6
thm:dichotomy cost-growth AR,E,H=cost-growth,LC,Q_fin,Qb,Qf,RG,S,S+ETH CC16, CC23, CC24
thm:dq-physical cost-growth AR,H=cost-growth DQ1, DQ2, DQ3, DQ4, DQ5, DQ6
thm:exact-certified-gap-iff-admissible cost-growth AR,E,H=cost-growth,Qb,Qf,RG,S,S+ETH CC17, CC18, CC20, IC31
thm:generalized-dominance cost-growth H=cost-growth HD9
thm:generalized-saturation-possible cost-growth H=cost-growth HD6, HD20
thm:growth-time cost-growth H=cost-growth,Q_fin IV1, IV2, IV3
@@
thm:orbital-cost cost-growth H=cost-growth,Q_fin AC3, AC4
thm:regime-coverage cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH CC14, CC51
thm:substrate-degradation cost-growth H=cost-growth,Q_fin SE1, SE2, SE3, SE4
thm:tax-conservation cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH HD5
thm:tractable cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH,STR,TA CC70, CC71, CC72, CC73
thm:typed-completeness-static cost-growth AR,E,H=cost-growth,LC,Qb,Qf,RG,S,S+ETH CC76
cor:finite-budget-no-exact-admissibility exp-lb-conditional H=exp-lb-conditional PBC8, PBC7
cor:finite-budget-threshold-impossibility exp-lb-conditional H=exp-lb-conditional PBC8, PBC7
cor:no-universal-survivor-no-succinct-bound exp-lb-conditional H=exp-lb-conditional PBC10
@@
prop:eventual-explicit-infeasibility exp-lb-conditional H=exp-lb-conditional PBC7
prop:least-divergence-point exp-lb-conditional H=exp-lb-conditional PBC5
prop:payoff-threshold exp-lb-conditional H=exp-lb-conditional PBC8, PBC9
prop:policy-closure-beyond-divergence exp-lb-conditional H=exp-lb-conditional PBC11, PBC12
@@
cor:query-obstruction-bool query-lb H=query-lb CC49, CC50
prop:checking-witnessing-duality query-lb H=query-lb WD5, WD3, WD4
prop:query-finite-state-generalization query-lb H=query-lb CC50
prop:query-randomized-robustness query-lb H=query-lb CC50
prop:query-randomized-weighted query-lb H=query-lb CC50
prop:query-tightness-full-scan query-lb H=query-lb CC50
prop:query-weighted-transfer query-lb H=query-lb CC50
@@
cor:practice-bounded tractable-structured H=tractable-structured CC81
cor:practice-structured tractable-structured H=tractable-structured CC82
cor:practice-symmetry tractable-structured H=tractable-structured CC72
cor:practice-tensor tractable-structured H=tractable-structured CC71
cor:practice-tree tractable-structured H=tractable-structured CC84
cor:practice-treewidth tractable-structured H=tractable-structured CC73, CC75
@@
prop:orbital-symmetry tractable-structured H=tractable-structured,Q_fin AC8
thm:complexity-dichotomy tractable-structured AR,E,H=tractable-structured,Q,S DC13, DC3
thm:six-subcases tractable-structured AR,E,H=tractable-structured,Q,S CC70, CC71, CC72, CC73, DC13
cor:ego-trap unspecified - (no derived Lean handle found)
cor:exact-no-competence-zero-certified unspecified AR IC41
cor:finite-state unspecified - BA2, BA4
cor:forced-finite-speed unspecified - BA10
cor:ground-state unspecified AR BA8
cor:ground-state-passive unspecified AR,Q_fin AC6
@@
cor:import-asymmetry unspecified AR SR4, SR5
@@
cor:phase-transition unspecified AR IE12, IE13
cor:physics-no-universal-exact-claim unspecified CR CC38
cor:posed-anchor-certified-gate unspecified - AQ8
cor:practice-unstructured unspecified - CC26
cor:rlff-abstain-no-certs unspecified AR IC30, IC40
cor:theorem-equilibrium unspecified AR IE3, IE4, IE5
cor:zero-gap unspecified AR GE4, GE5
prop:abstain-guess-self-signal unspecified AR IC46
prop:abstention-frontier unspecified RG CC39
@@
prop:certified-confidence-gate unspecified AR IC38, IC39
prop:comp-thermo-chain unspecified AR DS5, DS6
prop:decision-equivalence unspecified AR DE1, DE2, DE3, DE4
@@
prop:discrete-state-time unspecified AR DS1, DS2
prop:empty-sufficient-constant unspecified DM DP6
prop:evidence-admissibility-equivalence unspecified AR IC17, IC20, IC21
prop:exact-requires-evidence unspecified AR IC35, IC36
prop:fraction-defined-under-bound unspecified AR IC45
prop:heisenberg-strong-nontrivial-opt unspecified AR HS3, HS6, HS5
prop:identifiability-convergence unspecified ID CC19
prop:insufficiency-counterexample unspecified DM DP7, DP8
@@
prop:integrity-prerequisite unspecified AR IE17
prop:integrity-resource-bound unspecified AR,S,S+ETH CC30, IC24, IC26
@@
prop:lorentz-discrete unspecified AR DS3, DS4
prop:mdp-tractable unspecified FO CC50
@@
prop:no-evidence-zero-certified unspecified AR IC40
prop:no-fraction-without-bound unspecified AR IC44
@@
prop:physical-claim-transport unspecified AR CT6, CT3, CT5
prop:physics-no-universal-exact unspecified CR CC15
prop:pose-anchor-object unspecified - AQ1, AQ2, AQ3
prop:posed-anchor-typed-exact unspecified - AQ4, AQ5, AQ6, AQ7
prop:refinement-strengthens unspecified RG CC69
prop:retraction-evidence-integrity unspecified RG CC70
prop:retraction-no-evidence-violates unspecified RG CC71
@@
prop:self-confidence-not-certification unspecified AR IC47
prop:sequential-anchor-refinement unspecified PSPACE DC23
prop:sequential-anchor-tqbf-reduction unspecified PSPACE DC27, DC28, DC29
prop:sequential-bounded-horizon unspecified BH CC49
prop:sequential-static-relation unspecified DET,T=1 CC46
@@
prop:snapshot-process-typing unspecified RA CC48, CC56, CC3
prop:static-stochastic-strict unspecified P $\neq$ coNP DC1
prop:static-stochastic-transfer unspecified PD CC52
prop:steps-run-scalar unspecified AR IC42, IC43
prop:stochastic-anchor-refinement unspecified $P$ DC19
prop:stochastic-anchor-strict-reduction unspecified $P$ DC24, DC25, DC26
prop:stochastic-bounded-support unspecified BS CC59
prop:stochastic-potential-duality unspecified $P$ DC33, DC34, DC31
prop:stochastic-product-tractable unspecified PD CC60
prop:stochastic-sequential-bridge-fail unspecified CE CC56
prop:stochastic-sequential-strict unspecified P $\neq$ $P$ DC2
prop:temporal-equilibrium unspecified AR IE10, IE11
@@
prop:typed-physical-transport-requirement unspecified AR PS3, PS4
prop:under-resolution-collision unspecified DM PI6, PI7
@@
thm:assumption-necessity unspecified - AN4, AN3
thm:bayes-from-counting unspecified - BC4, BC3, BC1, BC2, BC5
thm:bayes-optimal unspecified - FN7, FN14, FN12
thm:boolean-primitive unspecified - BA4
thm:bounded-acquisition unspecified - BA2
thm:checking-duality unspecified - WD3, WD1, WD2
thm:claim-integrity-meta unspecified CR,DC CC22
thm:competence-access unspecified AR,E,Q,S IA5, IA6
thm:competence-capacity unspecified AR,Q_fin CH1, CH5
thm:config-reduction unspecified AR,E,RG,S,TA CR1
thm:cost-asymmetry-eth unspecified S+ETH CC12, HD14
thm:counting-gap unspecified - BA10
thm:counting-gap-intro unspecified - BA10
thm:discrete-acquisition unspecified - BA3
thm:ec2-derived unspecified - IA11, IA12, IA13, IA7, IA9
thm:ec3-derived unspecified - FPT10, FPT4, FPT5, FPT6, FPT8
thm:energy-information unspecified - EI1
thm:entropy-rank unspecified - IT4, IT3
thm:fi-coincide unspecified - FI7, FI3, FI6
thm:fisher-rank-srank unspecified - FS2, FS1
thm:gap-physical unspecified AR,E,Q,S IA4
thm:information-gap unspecified AR,E,Q,S IA3
thm:landauer-structure unspecified - FP15
thm:measure-prerequisite unspecified - MN1, MN10, MN2, MN11
thm:nontriviality-counting unspecified - FP3, FP2, FP7, FP4, FP1, FP6, FP5
thm:orbital-transition unspecified AR,Q_fin AC1, AC5
@@
thm:physical-incompleteness unspecified AR PI3, PI5, PI4
thm:quotient-universal unspecified - QT2, QT1, QT3
thm:rate-distortion-bridge unspecified - RS3, RS1, RS5, RS2, RD3, RD2, RD1, RS4
thm:resolution-sufficient unspecified - BA5
thm:second-law-counting unspecified - FP8, FP14, FP10, FP9, FP12, FP13, FP11
thm:srank-physical unspecified - SR1, BA7, BA6
thm:tax-grows unspecified LC HD26
thm:thermo-derived unspecified - BA9
thm:topology-motion unspecified AR,E,Q,S CC16, CC24
thm:tur-bridge unspecified - TUR6, TUR1, TUR2, TUR5
thm:universe-membership unspecified - (no derived Lean handle found)
thm:wasserstein-bridge unspecified - W3, W1, W2, W4
Auto summary: indexed 222 claims by hardness profile (cost-growth=56; exp-lb-conditional=11; query-lb=8; succinct-hard=6; tractable-structured=12; unspecified=129).